OpenStudy (anonymous):

how do i convert 11 pie over 4 radians to degrees

OpenStudy (ranga):

(pi) radians = 180 degrees.

OpenStudy (anonymous):

so would i multiply 180 by 11 then divide by 4?

OpenStudy (ranga):

To convert 11 pie over 4 radians to degrees just multiply by 180 / (pi)

OpenStudy (anonymous):

multiply what by 180?

OpenStudy (jdoe0001):

\(\bf Degrees = \cfrac{{\color{red}{ Radians}}\cdot 180}{\pi}\\ \quad \\ Degrees =\cfrac{{\color{red}{ \frac{11\pi}{4}}}\cdot 180}{\pi}\)

OpenStudy (ranga):

11(pi) / 4 * 180/(pi) = 11 * 180 / 4

OpenStudy (anonymous):

so is the answer 495?

OpenStudy (unklerhaukus):

yes